OpenStudy (leann01xxx):

The quadratic formula is used to solve an equation and the result is 2±√36/6. The solution set for the equation is A. {-2/3, 4/3} B. {1, 3} C. {2}

OpenStudy (e.mccormick):

Well, if you simplify that, what do you get?

OpenStudy (leann01xxx):

I have no idea how to simplify this..

OpenStudy (e.mccormick):

Well, what is the square root of 36?

OpenStudy (leann01xxx):

6 im pretty sure

OpenStudy (e.mccormick):

\(2\pm \dfrac{\sqrt{36}}{6}\) \(2\pm \dfrac{6}{6}\) OK, what is next?

OpenStudy (leann01xxx):

Uh im not really sure. Im really bad at math and failed it, so..

OpenStudy (e.mccormick):

I am assuming the first... but you might have meant: \(\dfrac{2\pm \sqrt{36}}{6}\) It is a little unclear which by how it is written. 2±√(36)/6 = \(2\pm \dfrac{\sqrt{36}}{6}\) (2±√36)/6 = \(\dfrac{2\pm \sqrt{36}}{6}\) In either case, the first thing is getting thr square root of 36, which is 6.

OpenStudy (leann01xxx):

Ok, I still don't understand how to get the answer.

OpenStudy (e.mccormick):

It is simplification. Order of operations. PEMDAS Parenthesis Exponents Multiplication and Division Addition and Subtraction Now, the square root is related to exponents, so pretty high up there.

OpenStudy (leann01xxx):

I don't get this..

OpenStudy (leann01xxx):

I got it. It was A.

OpenStudy (e.mccormick):

1+2*3 Because of the order of operations, I multiply first: 1+6 7 But if it was: (1+2)*3 The paren make the addition happen first: 3*3 9 The order changes the answer. That is why it is a key concept. All this question is saying is that you need to use the order of operations to simplify the result. So the only other issue is exactly how it is written, which changes it a lot.